LEP EXPERIMENTS COMMITTEE
Minutes of the 41st meeting,
14 November 1995


 


OPEN SESSION

1. LEP performance in 1995 and expected future performance

S. Myers reported on the very successful LEP operations in 1995. After initial difficulties with commissioning the new bunch-train optics, a successful scan of the Z0 had been carried out. A total luminosity of 40 pb-1 was delivered at the Z0 (including 20 pb-1 off-peak). The scan was taken with 4 x 3 bunches at energies of peak, +2 GeV and -2 GeV.

The machine energy had now been raised to 65-70 GeV, with a fast start-up and excellent performance. A peak luminosity, 2.6 x 1031 cm-2s-1 had been achieved and 3.8 pb-1 had been delivered at higher energies (by the end of the 1995 cycle, these figures were 3.4 x 1031 cm-2s-1 and 6.3 pb-1, respectively). A vertical beam-beam tune shift, xy = 0.05 and a small betatron coupling parameter (emittance ratio), k2 = 0.005 have been measured at 65 GeV. These figures are encouraging signs for the future performance of LEP 2.

Bunch trains have achieved performances as good as the pretzel scheme, and record high currents (up to 8.4 mA) have been delivered for physics.
Operation with 4 x 4 bunches is difficult, but is not required for LEP 2. The most likely configuration for LEP 2 is 4 x 2 bunches; the optimum
bunch spacing is to be discussed at the Chamonix workshop in January 1996.

Regarding future performance, LEP 2 will begin the 1996 cycle at about 161 GeV total energy and, by October 1996, should be able to operate up
to about 175 GeV, well above W+W- threshold. In June 1997 the maximum energy will be 186 GeV (Phase IIIb), and in June 1998 it will be
192 GeV (Phase IV). In order to achieve the luminosity goal of 500 pb-1 in 3 years, a peak luminosity of about 1032 cm-2s-1 will be required.

2. Status of the LEP 2 superconducting RF system

D. Boussard presented a progress report on the LEP 2 sc RF cavities. A total of 40 modules will have been accepted by the end of 1995, exactly according to the target defined at the start of the year. The main stream of the work at CERN on preparing the modules has become significantly more efficient, with high acceptance rates of cavities from industry and numerous improvements in the procedures for installation of couplers, testing, conditioning and cavity performance recovery. Sixteen sc RF modules (64 cavities) are presently installed at LEP, providing two thirds of the total RF voltage.

No major difficulties have been experienced so far with the sc cavities. However there have been difficulties with cavity "microphonics" (mechanical vibrations) which appeared with beam; they have been overcome by operating the cavities with a different de-tuning. Improvements in the RF system regulation loops will be implemented to cope with the larger impedance of the sc cavities (at the RF frequency) and the new mode of operation of the tuners. The performance and operational experience with the main RF couplers-which are protected against an earlier multipacting problem by DC-biassing-and with the HOM couplers are both excellent.

A major work programme to install an additional 18 modules during the 95/96 shut-down is about to begin, with a further 8 modules to be installed during the long shut-down in the middle of 1996. The total RF cavities installed at LEP by October 1996 will be 120 Cu and 168 sc (42 modules).

3. LEP energy calibration

J. Yamartino reported on the work of the LEP Energy Working Group for the 1995 Z0 scan. Numerous improvements were made relative to the earlier 1993 scan. These included energy calibrations at both the beginning and end of fills, improved monitoring of B fields with NMR probes in the tunnel magnets, conditioning (modulating) to improve magnet stability, and extensive logging of all parameters related to the beam energy. In order to control the effect of finite vertical dispersion at the interaction points, around 1000 vernier scans had been performed and the vertical offsets were controlled to a precision of about 0.2 mm. The estimated uncertainty in collision energy from this effect is below 0.5 MeV. A peculiar noise pattern of several MeV amplitude had been observed on the ring NMR probes, which disappeared each day between midnight and 4 a.m. It was eventually traced to a ground current return path from the Swiss and French trains through the LEP vacuum chamber! Although there is much work to be done, the conclusion is that the 1995 scan was a success and will result in improved precision of the Z0 mass and width.

4. Reports on the LEP experiments

All detectors reported very successful operation during 1995. A total of 4-4.5M Z0 hadronic decays have been collected per experiment during the entire operation of LEP 1. Each of the experiments congratulated the LEP machine team for an impressive transition to higher energies, with record luminosities and extremely low background conditions recorded at energies between 130 and 140 GeV. All observations at higher energy are in accord with the Standard Model, and charginos in the mass range up to 60-65 GeV are close to being excluded.

Concerning beam requests in 1996, all experiments support the proposed plan to start at a centre-of-mass energy about 161 GeV and then to operate near
175 GeV from October onwards, following the long shut-down to install more sc RF cavities. The preliminary requests for Z0 calibration data at the start
of 1996 are as follows: OPAL (1 pb-1), ALEPH (1-3 pb-1), DELPHI (3 pb-1), and L3 (4-6 pb-1). The large request from L3 reflects the requirements for precision alignment of the small-lever-arm inner tracking detector with back-to-back tracks, and of the outer muon spectrometer with m+m- events.

CLOSED SESSION

Present: W. Bartel, S. Bethke, W. Blum, M. Calvetti, T. Camporesi, J.D. Dowell (Chairman), L. Foà, B. Foster,
                 W. Hollik, K. Hübner, G. Kantardjian, J. Kirkby (Secretary), K.H. Kissler,  M. Klein, B. Mansoulié,
                 G. Matthiae, R. Mount, S. Myers, J. Panman,  L. Pape, I. Videau,  A. Weidberg, P. Wells, D.O. Williams
                 and F. Zwirner.

Apologies: P. Hansen

1. Approval of the minutes of the 40th meeting

The minutes of the 40th meeting (LEPC 95-6/LEPC 40) were approved without modification.

2. Chairman's report

The Chairman reported on the proposal to upgrade the energy of LEP 2 to Phase IV, which involves purchasing an extra 32 sc RF cavities to raise the centre-of-mass energy to 192.4 GeV [1]. The LEP 2 energy upgrade has been discussed and supported by the Research Board, SPC and CC. Before final approval can be given, however, the proposed financial plan must be agreed by the FC and Council.

It is proposed to fund the upgrade (36MCHF) from several sources, which include a reduction in LEP operation by an additional 2 weeks per year. Taken together with reduced operations due to the increased energy consumption at high energy, the Research Director indicated that the maximum yearly running period of LEP 2 from 1997 onwards would be 1 June - 31 October. The committee expressed its regret at this severe reduction planned for LEP running, which amounts to about 2 months per year, and urged CERN management to restore as much LEP 2 beam time as possible.

The decision on the Z0 scan was taken on 25 July by the LEPC Chairman and the Directors of Accelerators and Research, in consultation with the LEP Co-ordinator and the Spokesmen of the four LEP detectors. Although the four experiments had been evenly divided, a decision was made to proceed when it became clear that there would be no interference with the commissioning of higher beam energies. The scan achieved its objective of 20 pb-1 of calibrated off-peak data.

3. Discussion on aspects of the open session

The committee took great pleasure in congratulating the machine personnel for the excellent transition of LEP to higher energies, Ecm = 130-140 GeV, and for the successful commissioning of the superconducting RF system. The goal of 5 pb-1 at high energy would clearly be reached. The committee was impressed by how well the experiments have operated at high energy and by the rapid production of physics results.

The committee also congratulated the LEP machine teams and the Energy Working Group for successfully carrying out the Z0 scan under the challenging conditions of bunch-train operation and commissioning of the sc RF cavities. The new scan and the deepening understanding of the systematic errors on the LEP beam energy should result in increased precision and confidence in the final measurement of the Z0 mass and width from LEP.

4. Discussion on Z0 calibration data during LEP 2 operation.

The committee discussed the preliminary requests from the four LEP detectors for Z0 calibration data during high energy operation of LEP, which vary between 1 pb-1 (OPAL) and 4-6 pb-1 (L3). From the viewpoint of machine operation, Z0 calibration data during high-energy running presents no difficulties-the machine energy is already ramped through 45 GeV on its way to higher energy. Until the performance of LEP 2 is better known, however, it is difficult to reliably convert these calibration requests into days; pessimistically, however, the luminosity at the Z0 may be 0.3 pb-1 per day. In view of the short expected running time during 1996 (next item), the committee therefore proposed a guide-line of about 1 pb-1 Z0 calibration data at the start of the 1996 cycle, and encouraged the LEP experiments to explore how well their detectors could be calibrated with alternative sources that are present in the high-energy data (two-photon events, cosmics, etc.). This would be re-discussed at the next meeting in June.

5. Report from the LEP Co-ordinator; draft 1996 LEP schedule

The LEP Co-ordinator presented the draft LEP schedule for 1996 (attached). Following cold check-out, LEP will start up on 17 June 1996. After the Z0 detector calibration data, LEP will operate at W-pair threshold (~161 GeV) until 19 August. Further installation of sc RF cavities for LEP 2 Phase II will take place during a long shut-down of 7.5 weeks between 19 August and 9 October. LEP will operate the remainder of 1996 (9 October - 17 November) at about 175 GeV.

The committee discussed the physics objective of the run at 161 GeV-namely to measure the W mass. A measurement of the W mass that is competitive with the expected Tevatron measurement would require about 50 pb-1 per experiment, giving a combined LEP error of s (MW) = 100 MeV. This is considerably more than the integrated luminosity at this energy that is expected in the present schedule. If a competitive W mass measurement at threshold becomes a goal, it might suggest a re-optimization of the RF cavity installation schedule.

The total operating periods for LEP in 1996 are 72 days of physics, 20 days of MD and 10 days of set-up. This indicates a total of 100 days of LEP beams in 1996, to be compared with 168 days in 1995.

6. A.O.B.

In closing the meeting, the Chairman warmly thanked the outgoing members of the LEPC, W. Bartel and A. Weidberg.
He also thanked T. Camporesi for his excellent work as LEP Co-ordinator and welcomed the new LEP Co-ordinator, P. Wells,
who will start on 1 April 1996.

7. LEPC meetings in 1996

The dates of the next meeting of the LEPC will be Tuesday-Wednesday, 4-5 June 1996. The tentative dates of LEPC meetings in 1996 are:
 

In addition, there will be meeting in mid-September, with dates to be finalised shortly. This replaces the Cogne meeting of the LEPC to review LEP physics and plans, which has now been postponed until 1997.
  DOCUMENTS RECEIVED

[1]    S. Myers and C. Wyss, LEP 2 Energy Upgrade, CERN LEP2 Note 95-34, LEPC 95-11/M113.